OpenStudy (anonymous):

please help me i need to find θ - the question is sinθ=negative square root 3 /2 , θ ~[-π ,π] an explanation would be appreciated :)

OpenStudy (anonymous):

great -.- a smarter answer perhaps?

OpenStudy (anonymous):

So you have: \[\sinθ = -\frac{ \sqrt{3} }{ 2 }\] to find θ take the arcsin of both sides. \[\sin^{-1} (\sinθ) = \sin^{-1} -\frac{ \sqrt{3} }{ 2}\] You'd get: θ = -60 degrees or in radians: θ = \[-\frac{ π }{ 3 }\] Think this is what you were asking for.

OpenStudy (anonymous):

thankyou so much :)

OpenStudy (anonymous):

No problem :)